Horizon Telecom sold $300,000 worth of 120-day commercial paper for $298,000. What is the dollar amount of interest paid on the commercial paper? What is the effective 120-day rate on the paper?
Answer #1

Dollar amount of interest paid on the commerical paper=300000-298000=2000

Effective 120-day rate=(300000/298000)-1=0.6711%